OpenStudy (anonymous):

Is it possible to use product rule for the denominator then the quotient rule for the rest of this equation? Given that y=3x^2+6x-7/(x+1)^2 Show that dy/dx = 20/(x+1)^3

OpenStudy (unklerhaukus):

@hartnn

hartnn (hartnn):

y=(3x^2+6x-7)/(x+1)^2 =(3x^2+6x-7)/(x+1)^(-2) so, yes you can use product rule now. but complexity for both product and quotient rule will be almost same...

OpenStudy (anonymous):

Sorry I meant to say chain rule seperately for the denominator, then using the quotient rule

hartnn (hartnn):

chain rule seperately for the denominator <----no just quotient rule.

OpenStudy (anonymous):

So it would be 3x^2+6x-7/(x+1)^2/(2x+2) I have also tried product rule but can't seem to get the right answer

hartnn (hartnn):

no no, chain rule won't apply in denominator, it will apply in nemerator, when you will differentiate (x+2)^2 you know quotient rule, right ?

OpenStudy (anonymous):

Yes, when I left it as it is and just used quotient rule I get 6x^3+3x^2+13/(x+1)^4 This is after I have expanded the bracketts

hartnn (hartnn):

let me try the numerator (6x+6)(2x+2)^2 - 4(3x^2+6x-7) (2x+2) = 24 [(x+1)^3] - 8 (3x^3+6x^2-7x+3x^2+6x-7) = 24 [(x+1)^3] - 8 (3x^3 +9x^2-x-7) hmm...did i do any mistake anywhere ?

hartnn (hartnn):

because numerator is not coming 20.....

OpenStudy (anonymous):

Not sure this question was for the least amount of marks on the exam but seems to be the most difficult. Also what happened to the your denominator?

hartnn (hartnn):

in quotient rule Denominator is squared so, it will remain as (x+1)^4 from beginning

hartnn (hartnn):

so our numerator must come out to be 20 (x+1) let me proceed 24 [(x+1)^3] - 8 (3x^3 +9x^2-x-7) = 24 x^3 + 24*3x^2 +24*3x+24- 24x^3 -24*3x^2 +8x +56 = 24*3x +8x +80 = 80 x+ 80 = 80(x+1) O.o we are close

hartnn (hartnn):

ohh... i took, 2x+2 instead of (x+1) now i am getting 20

hartnn (hartnn):

(6x+6)(x+1)^2 - 2(3x^2+6x-7) (x+1) = 6 [(x+1)^3] - 2(3x^3+6x^2-7x+3x^2+6x-7) = 6 [(x+1)^3] - 2 (3x^3 +9x^2-x-7) =6 [(x+1)^3] - 2 (3x^3 +9x^2-x-7) = 6x^3 + 6*3x^2 +6*3x+6- 6x^3 -6*3x^2 +2x +14 =6*3x +2x +20 = 20x+20 = 20(x+1) it definitely wasn't easy...but just involved lot of algebra

OpenStudy (anonymous):

Ah ok so you then cancel out (x+1) from the numerator. Thank you very much, this has helped me a lot

hartnn (hartnn):

welcome ^_^